OpenStudy (anonymous):

Solve the systems of equations by graphing y = -3x + 18 3x - 2y = 0

OpenStudy (anonymous):

Can you help me step by step please?

OpenStudy (aravindg):

ok use the method told in Notes section in my tutorial http://openstudy.com/updates/50f8cc3fe4b027eb5d9a556a

OpenStudy (anonymous):

So I do not use the substitution or elimination method?

OpenStudy (aravindg):

Solve the systems of equations by graphing

OpenStudy (aravindg):

so why are you thinking of elimination and substitution ?

OpenStudy (anonymous):

Because i need to solve to get the points, then graph?

OpenStudy (anonymous):

subtract 3x from both sides of second y=-3x+18 -2y=-3x Divide each term in the equation by -2. y=-3x+18 y=(3x)/(2) Create a graph to locate the intersection of the equations. The intersection of the system of equations is the solution. y=-3x+18 y=(3x)/(2)

OpenStudy (anonymous):